OpenStudy (anonymous):

If f(x) = 2x - 2 and g(x) = 3x - 1, find (f - g)(4). Answer -17 5 -5 17

OpenStudy (anonymous):

-5

OpenStudy (jamesj):

Now why? Just giving the answer perhaps isn't very helpful. The reason is \[ (f-g)(x) = f(x) - g(x) \] Now that being the case, \[ (f-g)(4) = f(4) - g(4) \] Hence you just need to evaluate f(4) and g(4), and then subtract one value from the other.

OpenStudy (anonymous):

.-. what?

OpenStudy (jamesj):

You want to find (f - g)(4). Now what does that mean? It means the function f - g evaluated at 4. So what does f - g mean as a function? Well, the function f - g is defined as \[ (f-g)(x) = f(x) - g(x) \] Hence if you want to find (f-g)(4), then \[ (f-g)(4) = f(4) - g(4) \] Therefore to find the value of (f-g)(4), you need to find first 1. The value of f(4) 2. The value of g(4) 3. And then subtract them to find the value f(4) - g(4) So step 1: given that f(x) = 2x - 2, it follows that \[ f(4) = 2(4) - 2 = 8 - 2 = 6 \] Now you do step 2 and step 3.

OpenStudy (anonymous):

f(4) = 8-2=8-2=6 6=6

OpenStudy (jamesj):

Step 2. What is the value of g(4), given that g(x) = 3x - 1?

OpenStudy (anonymous):

g(4)= 3(4)-1 g(4)= 12 - 1 g(4)= 11

OpenStudy (jamesj):

Right. Now, step 3, the value of (f-g)(4) is f(4) - g(4) = .... what?